Find the margin of error if the confidence (significant) level is 95%.

1. n = 100, x = 80, s= 2.5

2. n=250, x = 100, s= 3.25

3. n= 300, x = 250, s = 4.15


1
Expert's answer
2022-04-06T07:57:40-0400

"E=\frac{z\sigma}{\sqrt{n}}"

1."E=1.96\times2.5/\sqrt{100}=0.49"

2. "E=1.96\times3.25/\sqrt{250}=0.40"

3. "E=1.96\times4.15/\sqrt{300}=0.47"